OpenStudy (anonymous):

3log2^5 = x and log2^5^3 = x are the same equation. true or false

OpenStudy (unklerhaukus):

\[3\log2^5 = x \] \[\log{2^5}^3 = x\] these equations are equivalent but they are not the same equation

OpenStudy (anonymous):

so false